Economic State Probability Return
Fast Growth 25% 24%
Slow growth 60% 9%
Recession 15% -6%

Compute the standard deviation of the expected return given these three economic states, their likilihoods, and the potential returns:
Answer #1

Expected return=respective return*respective probability

=(0.25*24)+(0.6*9)+(0.15*-6)=10.5%

probability Return probability*(return-Expected return)^2
0.25 24 0.25*(24-10.5)^2=45.5625
0.6 9 0.6*(9-10.5)^2=1.35
0.15 -6 0.15*(-6-10.5)^2=40.8375
Total=87.75%

Standard deviation=[Total probability*(return-Expected return)^2/Total probability]^(1/2)

which is equal to

=9.37%(Approx).